OpenStudy (anonymous):

Solve: 2x + 10 + 26 = 212 124 114 98 88

OpenStudy (anonymous):

2x + 10 + 26 = 212 what is 10+26=?

OpenStudy (anonymous):

let me do it surdawi i was here first haha jk

OpenStudy (anonymous):

Are we solving for x?

OpenStudy (anonymous):

@sue101 go ahead :P

OpenStudy (anonymous):

x = 88

OpenStudy (anonymous):

@SavannahSaturn don't just give the answer. Please show how you got it so the asker can learn.

OpenStudy (anonymous):

>:(

OpenStudy (anonymous):

@kymber fix your solution. 10+26=36

OpenStudy (anonymous):

2x + 10 + 26 = 212 2x + 36 = 212 2x + 36 - 36 = 212 - 36 2x = 176 2x/ 2 = 176/ 2 x = 88

OpenStudy (kymber):

 Woops.